the length of a rectangular field is two more than twice the width. the perimeter is 137 meters. find the dimensions.

Answer by addingup(3677)

2L + 2W = 137
L = 2W + 2 substitute for L above:
2(2W + 2) + 2W = 137
4W + 4 + 2W = 137
6W = 133
W = 22.167
and
L = 2W + 2 = 2(22.167) + 2 = 44.334 + 2 = 46.334
-------------------------------
Check:
2L + 2W; 2(46.334) + 2(22.167) = 92.668 + 44.334 = 137 Correct